# Client Selection for Federated Policy Optimization with Environment Heterogeneity

### v5

@@ Appendix B Proof of Lemma 6
Proof  By the definition of the Bellman operators, we have
$\left\|T^{\pi^{t+1}}_{I}\bar{V}^{t}-T_{I}\bar{V}^{t}\right\|$ $=\left\|\sum_{n=1}^{N}q_{n}\left(T^{\pi^{t+1}_{n}}_{I}\bar{V}^{t}-T_{n}\bar{V}^{t}\right)\right\|$
@@
$\left\|T^{\pi^{t+1}}_{I}\bar{V}^{t}-T_{I}\bar{V}^{t}\right\|$ $\leq\frac{2\gamma^{2}R_{\max}\kappa_{2}}{(1-\gamma)^{2}}+\frac{\gamma R_{\max}\kappa_{1}}{1-\gamma}+4\gamma\bar{\delta}+\bar{\epsilon}.$
@@ Appendix C Proof of Proposition 7
Proof  Given $\pi^{t+1}(a|s)=\sum_{m\in\mathcal{C}}q^{\prime}_{m}\pi^{t+1}_{m}(a|s),\forall s\in\mathcal{S},a\in\mathcal{A}$, we have
$\left\|T^{\pi^{t+1}}_{I}\bar{V}^{t}-T_{I}\bar{V}^{t}\right\|$ $=\left\|\sum_{m\in\mathcal{C}}q^{\prime}_{m}\sum_{n=1}^{N}q_{n}\left(T^{\pi^{t+1}_{m}}_{I}\bar{V}^{t}-T_{n}\bar{V}^{t}\right)\right\|$
@@
where step (a) follows from the contraction property of the Bellman operators and the triangle inequality. Step (b) follows from Lemma 13 and (7). Thus, by substituting (28) into the RHS of (27), we can conclude
@@
$\quad+2\gamma\sum_{m\in\mathcal{C}}q^{\prime}_{m}\delta_{m}+2\gamma\bar{\delta}+\sum_{m\in\mathcal{C}}q^{\prime}_{m}\epsilon_{m}.$
@@ Appendix E Proof of Proposition 9
Proof  Given that $\pi^{t+1}(a|s)=\sum_{m\in\mathcal{C}}q^{\prime}_{m}\pi^{t+1}_{m}(a|s),\forall s\in\mathcal{S},a\in\mathcal{A}$, we have
$\left\|T^{\pi^{t+1}}_{I}\bar{V}^{t}-T_{I}\bar{V}^{t}\right\|$ $=\left\|\sum_{m\in\mathcal{C}}q^{\prime}_{m}\sum_{n=1}^{N}q_{n}\left(T^{\pi^{t+1}_{m}}_{I}\bar{V}^{t}-T_{n}\bar{V}^{t}\right)\right\|$
@@
$\left\|T^{\pi^{t+1}_{m}}_{I}\bar{V}^{t}-T_{n}\bar{V}^{t}\right\|$ $\leq\left\|T^{\pi^{t+1}_{m}}_{I}\bar{V}^{t}-T^{\pi^{t+1}_{m}}_{m}\bar{V}^{t}\right\|+\left\|T^{\pi^{t+1}_{m}}_{m}\bar{V}^{t}-T_{m}\bar{V}^{t}\right\|+\left\|T_{m}\bar{V}^{t}-T_{n}\bar{V}^{t}\right\|$
@@
where the last inequality follows from Lemma 13 and (7). By substituting (35) into (34), we can conclude
$\leq\sum_{m\in\mathcal{C}}\sum_{n=1}^{N}q^{\prime}_{m}q_{n}\frac{\gamma R_{\max}\kappa_{m,n}}{1-\gamma}+\sum_{m\in\mathcal{C}}q^{\prime}_{m}\frac{\gamma R_{\max}\kappa_{m,I}}{1-\gamma}+\sum_{m\in\mathcal{C}}q^{\prime}_{m}\epsilon_{m}.$
@@ Appendix F Proof of Proposition 12
$\epsilon^{\prime}$ $=\acute{\epsilon}=\bar{\epsilon}.$
Thus, we can conclude

### v6

@@ Appendix B Proof of Lemma 6
Proof  By the triangle inequality, we have
$\left\|T^{\pi^{t+1}}_{I}V^{t}-T_{I}V^{t}\right\|$ $\leq\left\|T^{\pi^{t+1}}_{I}V^{t}-T^{\bar{\pi}^{t+1}}_{I}V^{t}\right\|+\left\|T^{\bar{\pi}^{t+1}}_{I}V^{t}-T^{\bar{\pi}^{t+1}}_{I}\bar{V}^{t}\right\|$
$\quad+\left\|T^{\bar{\pi}^{t+1}}_{I}\bar{V}^{t}-T_{I}\bar{V}^{t}\right\|+\left\|T_{I}\bar{V}^{t}-T_{I}V^{t}\right\|,$
from which we can obtain
$\left\|T^{\pi^{t+1}}_{I}V^{t}-T_{I}V^{t}\right\|$ $\leq\frac{\bar{\varepsilon}_{\theta}\sqrt{\left|\mathcal{A}\right|}R_{\max}}{1-\gamma}+2\gamma\bar{\varepsilon}_{w}+\left\|T^{\bar{\pi}^{t+1}}_{I}\bar{V}^{t}-T_{I}\bar{V}^{t}\right\|,$ (38)
by Lemma 16, the contraction property of the Bellman operators, and the definition of $\bar{\varepsilon}_{w}$.
To finish the proof, it suffices to bound the third term on the RHS of (38). By the definition of the Bellman operators, we have
$\left\|T^{\bar{\pi}^{t+1}}_{I}\bar{V}^{t}-T_{I}\bar{V}^{t}\right\|$ $=\left\|\sum_{n=1}^{N}q_{n}\left(T^{\pi^{t+1}_{n}}_{I}\bar{V}^{t}-T_{n}\bar{V}^{t}\right)\right\|$
@@
$\left\|T^{\pi^{t+1}}_{I}\bar{V}^{t}-T_{I}\bar{V}^{t}\right\|$ $\leq\frac{2\gamma^{2}R_{\max}\kappa_{2}}{(1-\gamma)^{2}}+\frac{\gamma R_{\max}\kappa_{1}}{1-\gamma}+4\gamma\bar{\delta}+\bar{\epsilon}.$ (43)
By substituting (43) into (38), we can obtain
$\left\|T^{\pi^{t+1}}_{I}V^{t}-T_{I}V^{t}\right\|$ $\leq\frac{\bar{\varepsilon}_{\theta}\sqrt{\left|\mathcal{A}\right|}R_{\max}}{1-\gamma}+2\gamma\bar{\varepsilon}_{w}+\frac{2\gamma^{2}R_{\max}\kappa_{2}}{(1-\gamma)^{2}}+\frac{\gamma R_{\max}\kappa_{1}}{1-\gamma}+4\gamma\bar{\delta}+\bar{\epsilon}.$
@@ Appendix C Proof of Proposition 7
Proof  Let $\tilde{\pi}^{t+1}(a|s)=\sum_{m\in\mathcal{C}}q^{\prime}_{m}\pi^{t+1}_{m}(a|s),\forall s\in\mathcal{S},a\in\mathcal{A}$ denote the expected output of a set $\mathcal{C}$ of local policies. By the triangle inequality, we have
$\left\|T^{\pi^{t+1}}_{I}V^{t}-T_{I}V^{t}\right\|$ $\leq\left\|T^{\pi^{t+1}}_{I}V^{t}-T^{\bar{\pi}^{t+1}}_{I}V^{t}\right\|+\left\|T^{\tilde{\pi}^{t+1}}_{I}V^{t}-T^{\tilde{\pi}^{t+1}}_{I}\bar{V}^{t}\right\|$
$\quad+\left\|T^{\tilde{\pi}^{t+1}}_{I}\bar{V}^{t}-T_{I}\bar{V}^{t}\right\|+\left\|T_{I}\bar{V}^{t}-T_{I}V^{t}\right\|,$
from which we can obtain
$\left\|T^{\pi^{t+1}}_{I}V^{t}-T_{I}V^{t}\right\|$ $\leq\frac{\tilde{\varepsilon}_{\theta}\sqrt{\left|\mathcal{A}\right|}R_{\max}}{1-\gamma}+2\gamma\bar{\varepsilon}_{w}+\left\|T^{\tilde{\pi}^{t+1}}_{I}\bar{V}^{t}-T_{I}\bar{V}^{t}\right\|,$ (44)
by Lemma 16, the contraction property of the Bellman operators, and the definition of $\bar{\varepsilon}_{w}$.
To finish the proof, it suffices to bound the third term on the RHS of (44). By the definition of the Bellman operators, we have
$\left\|T^{\tilde{\pi}^{t+1}}_{I}\bar{V}^{t}-T_{I}\bar{V}^{t}\right\|$ $=\left\|\sum_{m\in\mathcal{C}}q^{\prime}_{m}\sum_{n=1}^{N}q_{n}\left(T^{\pi^{t+1}_{m}}_{I}\bar{V}^{t}-T_{n}\bar{V}^{t}\right)\right\|$
@@
where step (a) follows from the contraction property of the Bellman operators and the triangle inequality. Step (b) follows from Lemma 15 and (7). Thus, by substituting (46) into the RHS of (45), we can obtain
@@
$\quad+2\gamma\sum_{m\in\mathcal{C}}q^{\prime}_{m}\delta_{m}+2\gamma\bar{\delta}+\sum_{m\in\mathcal{C}}q^{\prime}_{m}\epsilon_{m}.$ (47)
By substituting (47) into (44), we can conclude
$\left\|T^{\pi^{t+1}}_{I}V^{t}-T_{I}V^{t}\right\|$ $\leq\frac{\tilde{\varepsilon}_{\theta}\sqrt{\left|\mathcal{A}\right|}R_{\max}}{1-\gamma}+2\gamma\bar{\varepsilon}_{w}+\sum_{m\in\mathcal{C}}\sum_{n=1}^{N}q^{\prime}_{m}q_{n}\frac{\left(\gamma+\gamma^{2}\right)R_{\max}\kappa_{m,n}}{(1-\gamma)^{2}}$
$\quad+\sum_{m\in\mathcal{C}}q^{\prime}_{m}\frac{\gamma R_{\max}\kappa_{m,I}}{1-\gamma}+2\gamma\sum_{m\in\mathcal{C}}q^{\prime}_{m}\delta_{m}+2\gamma\bar{\delta}+\sum_{m\in\mathcal{C}}q^{\prime}_{m}\epsilon_{m}.$
@@ Appendix D Proof of Theorem 10
$\quad+\frac{\tilde{\varepsilon}_{\theta}\sqrt{\left|\mathcal{A}\right|}R_{\max}}{\left(1-\gamma\right)^{3}}+\frac{2\gamma\bar{\varepsilon}_{w}}{\left(1-\gamma\right)^{2}}$
@@ Appendix E Proof of Proposition 9
Proof  Let $\tilde{\pi}^{t+1}(a|s)=\sum_{m\in\mathcal{C}}q^{\prime}_{m}\pi^{t+1}_{m}(a|s),\forall s\in\mathcal{S},a\in\mathcal{A}$ denote the expected output of a set $\mathcal{C}$ of local policies. For any state $s$, we have
$\left|T_{I}^{\pi^{t+1}}V^{t}(s)-T_{I}V^{t}(s)\right|$ $\leq\left|T_{I}^{\pi^{t+1}}V^{t}(s)-T^{\bar{\pi}^{t+1}}_{I}V^{t}(s)\right|+\left|T^{\bar{\pi}^{t+1}}_{I}V^{t}(s)-T_{I}V^{t}(s)\right|.$ (53)
$\left|T_{I}^{\pi^{t+1}}V^{t}(s)-T^{\bar{\pi}^{t+1}}_{I}V^{t}(s)\right|$ $\leq\frac{\bar{\varepsilon}_{\theta}\sqrt{\left|\mathcal{A}\right|}R_{\max}}{1-\gamma}.$ (54)
To finish the proof, it suffices to bound the second term on the RHS of (53). By the definition of the Bellman operators, we have
$\left\|T^{\tilde{\pi}^{t+1}}_{I}V^{t}-T_{I}V^{t}\right\|$ $=\left\|\sum_{m\in\mathcal{C}}q^{\prime}_{m}\sum_{n=1}^{N}q_{n}\left(T^{\pi^{t+1}_{m}}_{I}V^{t}-T_{n}V^{t}\right)\right\|$
$\leq\sum_{m\in\mathcal{C}}q^{\prime}_{m}\sum_{n=1}^{N}q_{n}\left\|T^{\pi^{t+1}_{m}}_{I}V^{t}-T_{n}V^{t}\right\|.$ (55)
@@
$\left\|T^{\pi^{t+1}_{m}}_{I}V^{t}-T_{n}V^{t}\right\|$ $\leq\left\|T^{\pi^{t+1}_{m}}_{I}V^{t}-T^{\pi^{t+1}_{m}}_{m}V^{t}\right\|+\left\|T^{\pi^{t+1}_{m}}_{m}V^{t}-T_{m}V^{t}\right\|+\left\|T_{m}V^{t}-T_{n}V^{t}\right\|$
@@
where the last inequality follows from Lemma 15 and (7). By substituting (56) into (55), we can obtain
$\left\|T^{\pi^{t+1}}_{I}V^{\pi^{t}}-T_{I}V^{\pi^{t}}\right\|$ $\leq\sum_{m\in\mathcal{C}}\sum_{n=1}^{N}q^{\prime}_{m}q_{n}\left\|T^{\pi^{t+1}_{m}}_{I}V^{\pi^{t}}-T_{n}V^{\pi^{t}}\right\|$
$\leq\sum_{m\in\mathcal{C}}\sum_{n=1}^{N}q^{\prime}_{m}q_{n}\frac{\gamma R_{\max}\kappa_{m,n}}{1-\gamma}+\sum_{m\in\mathcal{C}}q^{\prime}_{m}\frac{\gamma R_{\max}\kappa_{m,I}}{1-\gamma}+\sum_{m\in\mathcal{C}}q^{\prime}_{m}\epsilon_{m}.$ (57)
By substituting (57) into (53), we can conclude
$\left\|T^{\pi^{t+1}}_{I}V^{t}-T_{I}V^{t}\right\|$ $\leq\frac{\tilde{\varepsilon}_{\theta}\sqrt{\left|\mathcal{A}\right|}R_{\max}}{1-\gamma}+\sum_{m\in\mathcal{C}}\sum_{n=1}^{N}q^{\prime}_{m}q_{n}\frac{\gamma R_{\max}\kappa_{m,n}}{1-\gamma}$
$\quad+\sum_{m\in\mathcal{C}}q^{\prime}_{m}\frac{\gamma R_{\max}\kappa_{m,I}}{1-\gamma}+\sum_{m\in\mathcal{C}}q^{\prime}_{m}\epsilon_{m}.$
@@ Appendix F Proof of Proposition 14
$\epsilon^{\prime}$ $=\acute{\epsilon}=\bar{\epsilon}+\frac{\hat{\varepsilon}_{\theta}\sqrt{\left|\mathcal{A}\right|}R_{\max}}{1-\gamma},$
where $\hat{\varepsilon}_{\theta}$ is equal to $\bar{\varepsilon}_{\theta}$ and $\tilde{\varepsilon}_{\theta}$ for full participation and partial participation, respectively. Thus, we can conclude